HOW DO I COMPLETE THE NUMBER LINE:
-82 ___ ___ ___ ___ -70
THERE ARE 4 SPACES BETWEEN -82 AND -70

Found 2 solutions by ewatrrr, stanbon:
Answer by ewatrrr(24785)

 
Hi,
-82 ___ ___ ___ ___ -70 Note: -70-(-82) = 12, 12/5 = 2.4
Add 2.4 for each new space
-82, -79.6, etc...

Answer by stanbon(75887)

HOW DO I COMPLETE THE NUMBER LINE:
-82 ___ ___ ___ ___ -70
THERE ARE 4 NUMBERS BETWEEN -82 AND -70
-----
Range: 12
# of gaps: 5
----
width of gaps 12/5 = 2 2/5
-----
Left limit number: -82
1st number from left -82+ (2 2/5) = -79 3/5
2nd number from left -79 3/5 + (2 2/5) = -77 1/5
3rd number from left -77 1/5 + 2 2/5 = -74 4/5
4th number from left -74 4/5 + (2 2/5) = -72 2/5
Right limit number: -72 2/5 + (2 2/5) = -70
===================================================
Cheers,
Stan H.